OpenStudy (anonymous):

find the midpoint of the segment that has endpoints at (-4,-5) and (10, -2)

OpenStudy (anonymous):

take the average in each coordinate

OpenStudy (anonymous):

the average of (-4\) and \(10\) is \(\frac{-4+10}{2}=\frac{6}{2}=3\) so that is the first coordinate of the midpoint second is found the same way, but using the \(y\) coordinates